The product of X and 250 is the same as the sum of 12X and 6188. Find the Number.

Answer by london maths tutor(243)

The product of X and 250 = 250*X
The sum of 12X and 6188 = 12X + 6188
Therefore, 250X = 12X + 6188
250X - 12X = 6188
238X = 6188
X = 6188/238
X = 26